OpenStudy (anonymous):

What is the simplified form of (x)-(y^5)/(x^2)/(y^2)? I'll draw it out for you

OpenStudy (anonymous):



OpenStudy (zzr0ck3r):

= x/y^4 * y^2/x^2 = (x/x^2)(y^2/y^4) = (1/x)*(1/y^2) = x^(-1) * y^(-2)

OpenStudy (anonymous):

^^^ It's cannot be negative exponents... I'll draw the solutions for youu.

OpenStudy (anonymous):